40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 64 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  13890   Thu May 24 20:31:03 2018 gautamConfigurationALSDFD noises

A couple of months ago, I took 21 measurements of the delay line transfer function. As shown in Attachment #2, the unwrapped phase is more consistent with a cable length closer to 45m rather than 50m (assuming speed of light is 0.75c in the cable, as the datasheet says it is).

Attachment #1 shows the TF magnitude for the same measurements. There are some ripples consistent with reflections, so something in this system is not impedance matched. I believe I used the same power splitter to split the RF source between delayed and undelayed paths to make these TFs as is used in the current DFD setup to split the RF beatnote.

Quote:
 

I had made some TF measurements of the delay sometime ago, need to dig up the data and see what number that measurement yields.

  13889   Thu May 24 19:41:28 2018 gautamConfigurationALSBeathMouth reinstalled on PSL table

Summary:

  • DC light power incident on beat PD is ~400uW from the PSL and ~300uW from EX.
  • These numbers are consistent with measured mating sleeve and fiber coupler losses.
  • However, I measure an RF beatnote of 80mVpp (= -18dBm). This corresponds to a mode matching efficiency of ~15%, assuming InGaAs efficiency of 0.65A/W.

I find this hard to believe.

Details:

  • I took this opportunity to clean the fiber tips on the PSL table going into the BeatMouth.
  • PSL light power going into the BeatMouth is 2.6mW. Of which ~400uW reaches the Beat PD (measured using my new front panel monitor port).
  • Similarly, 1mW of EX light reaches the PSL table, of which ~300uW reaches the Beat PD.
  • The RF amplifier gain is 20dB, and RF transimpedance is 50 ohms.
  • Using the (electrical) 20dB coupled port on the front panel, I measured a beat signal with 8mVpp. So the actual beat note signal is 80mVpp.

Discussion:

As I see it, the possibilities are:

  1. My measurement technique/calculation is wrong.
  2. The beat PD is broken has optoelectronic different that is significantly different from specifications.
  3. The non-PM fiber lengths inside the beat box result in ~15% overlap between the PSL and EX beams. Morever, there is insignificant variation in the electrical beat amplitude as monitored on the control room analyzer. So there is negligible change in the polarization state inside the BeatMouth.

I guess #3 can be tested by varying the polarization content of one of the input beams through 90 degrees.

  13888   Thu May 24 16:08:12 2018 KiraUpdatePEMparts list

We will need to order a few things for our final setup.

  1. 1U box to place the heater circuit and temperature circuits in. The minimum depth that will fit all the electronics is 10 inches according to my sketch.
    • I found two possibilities online for this. I don't know exactly what our budget is, but this one is $144. According to the datasheet, the front panel is less than 19 inches wide, so if we are to order this one, I've adjusted the width of the panel I designed to match the width of the panel that comes with it I've labeled it in the attached file as 1U-panel-1.
    • The other possibility is this one. It comes with handles already which is quite nice. I wasn't able to find a price for it on the website.
  2. Front panel for the 1U box and block panel. I've attached them as .fpd files below in one .zip file. Not sure if this is the correct way to attach them, though.
  3. We'll also need a 16 gauge cable that has 6 wires bunched together. This is to connect up the heater circuit and AD590s. The other cables that we will need can be found in the lab.
  13887   Thu May 24 15:18:12 2018 KiraUpdatePEMPID loop restarted

Rana said that it wasn't necessary to gather more data on the temperature fluctuations so I have reconnected the heater circuit and restarted the PID loop with the can on the seismometer.

  13886   Thu May 24 13:06:17 2018 gautamConfigurationALSDFD noises

Summary:

  1. The DFD noise floor is ~0.5Hz/rtHz at 100Hz (see Attachment #2).
  2. I cannot account for the measured noise floor of the DFD system. The Marconi noise and the AA noise contributions should be neglibible at 100Hz.
  3. This SURF report would lead me to believe that the delay line cable length is 50m. But my calibration suggests it is shorter, more like 40m (see Attachment #1). I had made some TF measurements of the delay sometime ago, need to dig up the data and see what number that measurement yields.

Details and discussion: (diagrams to follow)

  • Delay line linearity was checked by driving the input with Marconi, waiting for any transient to die down, and averaging the I and Q inputs to the phase tracker servo (after correcting for the daughter board TF) for 10 seconds. The deg/MHz response was then calculated by trigonometry. Not sure what to make of the structure in the residuals, need to think about it.
  • DFD noise was checked by driving the DFD input with a Marconi at 50MHz, RF level = 8dBm, which are expected parameters for nominal ALS operation. In this configuration, I measured the spectrum of the phase tracker output. I then used the calibration factor from the above bullet to convert to Hz/rtHz.
  • The electronics noise contribution of the daughter board was calibrated to Hz/rtHz by using the Marconi to drive the DFD input with a known FM signal (mod depth ~0.05), and using the SR785 to measure the power of the FM peak. This allows one to back out the V/Hz calibration constant of the delay line. I tweaked the carrier frequency until the ratio of power in I channel to Q channel (or the other way around) was >20dB before making this measurement.
  • I have no proof - but I suspect that the whole host of harmonics in the noise spectrum is because the 1U AA chassis sits directly on top of some Sorensen power supplies. These Sorensens power the frequency distribution box in the LSC rack, so the simplest test to confirm would be to turn off the RF chain, and then Sorensens, and see if the peaky features persist.
  13885   Thu May 24 10:16:29 2018 gautamUpdateGeneralAll models on c1lsc frontend crashed

All models on the c1lsc front end were dead. Looking at slow trend data, looks like this happened ~6hours ago. I rebooted c1lsc and now all models are back up and running to their "nominal state".

  13884   Wed May 23 19:24:37 2018 Udit KhandelwalSummaryGeneralSummary 05/23/2018

Tip-Tilt Redesign Project with Koji:

Did further itirations to the ECD backplate. Going to determine minimum thickness between magnet hole and plus sign for eddy current damping.

Chamber optical table layouts

Finished the positioning of optics and instruments in SolidWorks for the Vertex chambers. The reference for positioning is "40m_upgrade_layout_Dec2012.dwg", and solidworks files I created are in the main 40m CAD folder.

  13883   Wed May 23 17:58:48 2018 gautamUpdateIOOMC1 Coil Driver pulled out
  • Marked up schematic + photo post changes uploaded to DCC page.
  • There was a capacitor in the DAQ monitor path making a 8kHz corner that I now removed (since the main point of this front panel HPF monitor point is to facilitate easy coil driver noise debugging, and I wanted to be able to use the SR785 out to high frequencies without accounting for an additional low pass). Transfer function from front panel LEMO input to front panel LEMO monitor is shown in Attachment #1.
  • Voltage noise measured at DB25 output (with the help of a breakout cable and SR560 G=100) with front panel LEMO input terminated to 50ohm, Bias input grounded, and pin1 of U21A grounded (i.e. watchdog enabled state) is shown in Attachment #2. This measurement was taken on the electronics bench.
  • Inside the lab (i.e. coil driver board plugged into eurocrate), the noise measured in the same way looks identical to what was measured in elog13870.
  • I tried repeating the measurement by powering the board using an bench power supply and grounding the bias input voltage near 1X6, and the strange noise profile persists. So this supports the hypothesis that some kind of environmental pickup is causing this noise profile. Needs more investigation. 

In any case, if it is indeed true that the optic sees this current noise, the place to make the measurement is probably the Sat. Box. Who knows what the pickup is over the ~15m of cable from 1X6 to the optic.

Quote:

Detailed report + photos to follow

 

  13882   Wed May 23 14:50:33 2018 KiraUpdatePEMtest setup with seismometer

This time the test went without issue. The first attachment is the data for the past 24 hours and the second attachment is the full data over 6 days. The average temperature fluctuations (from highest point to lowest point) for the can on was 0.43 C and for the can off it came out to 0.55 C. In addition the seismometer with the can off is about 1 C cooler than with the can on. I'd like to leave the can off until the end of the week so we can get a comparable data set for both the can on and off. Eventually I'll need to figure out a way to clamp the can down to the block in order to get better insulation and hopefully get even smaller temperature fluctuations.

  13881   Wed May 23 00:45:18 2018 johannesConfigurationGeneralAS port laser injection

I was planning to set up the additions to the AS table that are outlined in Attachment #1. Unfortunately the beam is too large for the 2mm clear aperture Faraday rotators that we have available at that position. I checked the 40m and QIL and found 5 Faraday isolators/rotators for 1064 nm total, but none have large enough aperture for the current setup. Some options for buying a larger aperture isolator are:

I wanted to leave the rest of the setup undisturbed at first, but I think a much easier solution would be to move the 2" focusing lens up by about 12", which moves the beam focus away from AS55 to where the Faraday will be placed, but we can re-focus it with another lens. I may have to change the mode-matching for the aux laser fiber slightly to accomodate this change, but if there are no other concerns I would like to start this work tomorrow (Wednesday).

  13880   Tue May 22 23:28:01 2018 gautamUpdateIOOMC1 Coil Driver pulled out

This work is now complete. MC1 coil driver board has been reinstalled, local damping of MC1 restored, and IMC has been locked. Detailed report + photos to follow, but measurement of the noise (for one channel) on the electronics workbench shows a broadband noise level of 5nV/rtHz (yes) around 100Hz, which is lower than what was measured here and consistent with what we expect from LISO modeling (with fast input terminated with 50ohm, slow input grounded).

Quote:

I have pulled out MC1 coil driver board from its Eurocrate, so IMC is unavailable until further notice.

 

  13879   Tue May 22 17:29:27 2018 keerthanaUpdateelogMEDM Diagram for the auxilary laser system control and display.

(keerthana, gautam, jon)

In the morning, Jon gave me an overview of the Auxiliary laser system which we are planning to setup. Based on the diagram he uploaded in the elog, I have made the MEDM diagram for controlling and displaying the parameters. Here the parameters which we will be controlling are temperature (in terms of voltage), oscilator frequency ( with the help of IFR 2023B), the frequency offset and the PID controls. The display includes the beat frequency, error signal voltage, control voltage and a switch to give feed back to the AUX laser. As the frequency counter is not connected at the moment, I haven't included its channel number in it. The screenshot of the diagram is attached with this. I am also considering to give a PID feedback to the slow control from the AUX feedback signal. The screen can be accessed from the PSL dropdown menu in sitemap.

  13878   Tue May 22 17:26:25 2018 gautamUpdateIOOMC1 Coil Driver pulled out

I have pulled out MC1 coil driver board from its Eurocrate, so IMC is unavailable until further notice. Plans:

  1. Thick film --> Thin Film
  2. AD797 --> Op27
  3. Remove Pots in analog actuation path.
  4. Measure noise
  5. Route HPF signal (UL DAQ Mon) to front panel. I think we should use the SMA connectors. That way, we have DC and AC voltage monitors available for debugging.

If there are no objections, I will execute Step #5 in the next couple of hours. I'm going to start with Steps 1-4.

  13877   Tue May 22 14:49:03 2018 KiraUpdatePEMtest setup with seismometer

It appears that one of the wires was disconnected overnight or this morning so I wasn't able to gather data over a full 24 hour period. Perhaps someone accidentally kicked it. I placed some cones in that area so hopefully the wires won't be in the way as much and I can get the data tomorrow. From the data I do have it seems that the seismometer is at a colder temperature when the can is not on, though it is difficult to see by how many degrees the temperature fluctuates. I've included the data from 5 days back to see the comparison.

  13876   Tue May 22 10:14:39 2018 Jon RichardsonConfigurationElectronicsDocumentation & Schematics for AUX-PSL PLL

 

Quote:

[Jon, Gautam]

Attached is supporting documentation for the AUX-PSL PLL electronics installed in the lower PSL shelf, as referenced in #13845.

Some initial loop measurements by Gautam and Koji (#13848) compare the performance of the LB1005 vs. an SR560 as the controller, and find the LB1005 to be advantageous (a higher UGF and phase margin). I have some additional measurements which I'll post separately.

Loop Design

Pickoffs of the AUX and PSL beams are routed onto a broadband-sensitive New Focus 1811 PD. The AUX laser temperature is tuned to place the optical beat note of the two fields near 50 MHz. The RF beat note is sensed by the AC-coupled PD channel, amplified, and mixed-down with a 50 MHz RF source to obtain a DC error signal. The down-converted term is isolated via a 1.9-MHz low-pass filter in parallel with a 50 Ohm resistor and fed into a Newport LB1005 proportional-integral (PI) servo controller. Controller settings are documented in the below schematic. The resulting control signal is fed back into the fast PZT actuator input of the AUX laser.

Schematic diagram of the PLL.

 

 

 

 

 

 

 

 

 

 

Hardware Photos

Optical layout on the PSL table.

 

PLL electronics installed in the lower PSL shelf.

 

Close-up view of the phase detector electronics.

 

Slow temp. (left) and fast PZT signals into the AUX controller.

 

AUX-PSL beat note locked at 50 MHz offset, from the control room.

 

 

  13875   Mon May 21 18:02:55 2018 keerthanaUpdateGeneralTesting of the new mini-circuits frequency counter

Today, I tested the new mini-circuit frequency counter by connecting it with the beat signal output. The frequency counter works fine. Now I am trying to get a display of the frequency in the computer screen using python programming. I have made the code for remotely changing oscilator frequency and it is saved in the folder 'ksnair'. A picture of the new mini circuits frequency counter is attached below. Part no: UFC-6000, S/N: 11501040012, Run: M075270.

  13874   Mon May 21 17:36:00 2018 poojaUpdateCamerasGigE camera image of ETMX

Today Steve and I tried to to capture the image of scattering of light by dust particles on the surface of ETMX using GigE camera. The image ( at gain =100, exposure time = 125000) obtained has been attached. Unlike the previous images, a creepy shape of bright spots was seen. Gautam helped us lock infrared light and see the image. A similar less intense shape was seen. This may be because of the dust on the lens.

  13873   Mon May 21 15:34:19 2018 gautamConfigurationElectronicsChannel hijacking history

Since we've been hijacking channels like there is no tomorrow for the AUX-PLL setup, I'm documenting the channel names here. The next time c1psl requires a reboot, I'll rename these channels to something more sensible. To find the channel mapping, Koji suggested I use this. Has worked well for us so far... We've labelled all pairs of wires pulled out of the cross connects and insulation taped the stripped ends, in case we ever need to go back to the original config.

Previously unused C1PSL Channels now used for AUX PLL

Channel name AI/AO Function
C1:PSL-126MOPA_126CURADJ AO Slow temperature control
C1:PSL-FSS_RFADJ AO Servo trigger TTL
C1:PSL-126MOPA_126PWR AI PLL error signal monitor
C1:PSL-126MOPA_DMON AI PLL control signal monitor
C1:PSL-FSS_PHCON AO

To mitigate integrator railing

  13872   Mon May 21 14:17:28 2018 KiraUpdatePEMtest setup with seismometer

I have attached the graph for the seismometer temperature fluctuations over 3 days. As we can see, there is a noticeable fluctuation in daily temperature as well as a difference between days in the maximum and minimum temperatures. I will repeat this test but take the can off to see if there's any difference between having the can on or off.

  13871   Mon May 21 10:15:35 2018 gautamUpdatePEMtest setup with seismometer

I guess it's fine for now while we are still finalizing the setup at EX, but we should eventually line up the seismometer axes with the IFO axes. Is there a photo of the orientation of the seismometer pre heater can tests? If not, probably good to make some sort of markings on the granite slab / seismometer to allow easy lining up of these axes...

  13870   Sun May 20 23:43:50 2018 gautamUpdateIOOCoil driver noise re-measurement

Summary:

In the IMC actuation chain, it looks like the MC1/MC3 de-whitening boards, and also all three MC optics' coil driver boards, are showing higher noise than expected from LISO modeling. One possible candidate is thick film resistors on the coil driver boards. The plan is to debug these further by pulling the board out of the Eurocrate and investigating on the electronics bench.

Why bother? Mainly because I want to see how good the IR ALS noise is, and currently, the PSL frequency noise is causing the measurement to be worse than references taken from previous known good times.

Details:

Sometime ago, rana suggested to me that I should do this measurement more systematically.

  • Attachments #1, #2 and #3 show noise measurements in various conditions for MC1, MC2 and MC3 respectively.
  • In the above three attachments, I stitched together multiple spans from the SR785, and so the bin-width is different. The data is downloaded from the analyzer normalized by the bin-width (PSD units).
  • The roll-off at ~800Hz in the orange trace for MC1 and MC3 is consistent with an 800 Hz LPF that was used for anti-image filtering from the old 2 kHz era.
  • While it may look like the analog de-whitening isn't being switched on in some of these plots, I confirmed by transfer function measurement that it is indeed switching.
  • Data used to make these plots are in Attachment #4. Unfortunately, the code requires some of my personal plotting librariesno and so I'm not uploading it.
  • Sketch of measurement setup is shown in Attachment #5. It is not indicated in the schematic, but the SR560 was operated in battery mode for this measurement.
  • For MC1, I did the additional measurement of terminating the LEMO input to the coil driver and measuring (what should have been) just the coil driver electronics noise. But this measurement doesn't look very clean, and hence, the decision to pull the board out to continue debugging.
  • While at 1X6, Rana tightened the LEMO connectors going to MC1. We should opportunistically do MC2 and MC3 as well.
  • Some changes to be made:
    • Thick film ---> thin film.
    • Reroute HPF-ed back-plane Vmon output to the front panel LEMO.

I've now restored all the wiring at 1X6 to their state before this work.

  13869   Sun May 20 17:43:01 2018 ranaUpdateElectronicsHow to choose resistors

Article from EE Times, describing why metal foil (NOT metal film) resistors are really better than wirewound when it comes to everything except high power dissipation.

Need to do some diggin to see if we can find ~1k metal foil resistors which can handle ~1W of heat.

Steve: here it is

  13868   Fri May 18 20:03:14 2018 PoojaUpdateCamerasTelescopic lens solution for GigE

Aim: To find telescopic lens solution to image test mass onto the sensor of GigE camera.

I wrote a python code to find an appropriate combination of lenses to focus the optic onto the camera keeping in mind practical constraints like distance of GigE camera from the optic ~ 1m and distance between the lenses need to be in accordance with the Thorlab lens tubes available. We have to image both the enire optic of size 3" and beam spot of 1" using this combination of lens. The image size that efficiently utilizes the entire sensor array is 1/4". Therefore the magnification required for imaging the entire optic is 1/12 and that for the beam spot is 1/4.

I checked the website of Thorlabs to get the available focal lengths of 2" lenses (instead of 1" lenses to collect sufficient power). I have tried several combination of lenses and the ones I found close enough to what is required have been listed below along with thier colorbar plots.

a) 150mm-150mm (Attachment 2 & 3)

With this combination, object distance varies like 50cm for 1" beam spot to 105cm for 3" spot. Therefore, it posses a difficulty that there is a difference of ~48cm in the distances between the optic and camera in the two cases: imaging the entire optic and the beam spot.

b) 125mm-150mm (Attachment 4 & 5)

With this combination, object distance varies like 45cm for 1" beam spot to 95cm for 3" spot. There is a difference of ~43cm in the distances between the optic and camera in the two cases: imaging the entire optic and the beam spot.

c) 125mm-125mm (Attachment 6 & 7)

The object distance varies like 45cm for 1" beam spot to 90cm for 3" spot. There is a difference of ~39cm in the distances between the optic and camera in the two cases: imaging the entire optic and the beam spot.

Sensitivity check was also done for these combination of lenses. An error of 1cm in object distance and 5mm in the distance between the lenses gives an error in magnification <2%.

The schematic of the telescopic lens system has been given in Attachment 8.

 

  13867   Fri May 18 19:59:55 2018 Jon RichardsonConfigurationElectronicsAUX-PSL PLL Characterization Measurements

Below is analysis of measurements I had taken of the AUX-PSL PLL using an SR560 as the servo controller (1 Hz single-pole low-pass, gain varied 100-500). The resulting transfer function is in good agreement with that found by Gautam and Koji (#13848). The optimal gain is found to be 200, which places the UGF at 15 kHz with a 45 deg phase margin.

For now I have reverted the PLL to use the SR560 instead of the LB1005. The issue with the LB1005 is that the TTL input for remote control only "freezes" the integrator, but does not actually reset it. This is fine if the lock is disabled in a controlled way (i.e., via the medm interface). However, if the lock is lost uncontrollably, the integrator is stuck in a garbage state that prevents re-locking. The only way to reset this integrator is to manually flip a switch on the controller box (no remote reset). Rana suggests we might be able to find a workaround using a remote-controlled relay before the controller.

  13866   Fri May 18 19:10:48 2018 keerthanaUpdateGeneralCode for adjusting the oscillator frequency remotly

Target: Phase locking can be acheived by giving a scan to the oscilator frequency. This frequency is now controlled using the knobe on the AM/FM signal generator 2023B. But we need to control it remotely by giving the inputs of start frequency, end frequency and the steps.

The frequency oscilator and the computer is connected with the help of GPIB Ethernet converter. The IP address of the converter I used is '192.168.113.109' and its GPIB address is 10.

I could change the oscilator frequency by changing the input frequency with the help of the code I made (Inorder to check this code, I have changed the oscilator frequency multiple times. I hope it didn't create trouble to anyone). Now I am trying to make this code better by adding certain features like numpy, argument parse etc, which I will be able complete by next week. I am also considering to develop the code to have a sliding system to control the oscillatory frequency.

For record: The maximum limit of frequency which i changed upto is 100MHz.

 

  13865   Fri May 18 18:14:18 2018 Udit KhandelwalSummaryGeneralSummary 05/18/2018

Tip-Tilt Suspension Design:

Designed a new ECD plate and changed dimensions of the side arms after discussing with Koji. After getting feedback on the changes, I will finish the assembly and send it to him to get approved for manufacturing.

 

  13864   Fri May 18 14:33:34 2018 KiraUpdatePEMtest setup with seismometer

Here is the result of my test. I think I'll leave the can on over the weekend because there's a long period of time where the seismometer heated up by 0.8 degrees so I can't fully see the fluctuations over a full 24 hour period.

  13863   Fri May 18 14:18:03 2018 gautamConfigurationElectronicsBasic MEDM Control Screen setup

I setup a basic MEDM screen for remote control of the PLL.

The Slow control voltage slider allows the frequency of the laser to be moved around via the front panel slow control BNC.

The TTL signal slider provides 0/5V to allow triggering of the servo. Eventually this functionality will be transferred to the buttons (which do not work for now).

The screen can be accessed from the PSL dropdown menu in sitemap. We can make this better eventually, but this should suffice for initial setup.

  13862   Fri May 18 09:13:41 2018 PoojaUpdateSUSColored GigE image

To obtain a colored version with good contrast of the grayscale image of scattering of light by dust particles on the surface of test mass, got using GigE camera. The original and colored images are attached here.

 

  13861   Fri May 18 07:41:01 2018 steveUpdateSUSclipping ITMX oplev

The ITMX oplev still clipping

Quote:

The ITMX oplev beam is clipping. It will be corected with locked arm

 

 

  13860   Thu May 17 18:05:01 2018 gautamUpdateSUSSR785 near 1X5

I'm working near 1X5 and there is an SR785 adjacent to the electronics rack with some cabling running along the floor. I plan to continue in the evening so please leave the setup as is.

During the course of this work, I noticed the +15V Sorensen in 1X6 has 6.8 A of current draw, while Steve's February2018 label says the current draw is 8.6A. Is this just a typo?

Steve: It was most likely my mistake. Tag is corrected to 6.8A


I'm still in the process of electronics characterization, so the SR785 is still hooked up. MC3 coil driver signal is broken out to measure the output voltage going to the coil (via Gainx100 SR560 Preamp), but MC is locked.

  13859   Thu May 17 15:38:19 2018 KiraUpdatePEMtest setup with seismometer

I've moved my setup to the actual seismometer. I attached the temperature sensor to the seismometer (attachment 1) with duct tape, though this is temporary. I will be monitoring the temperature fluctuations of the seismometer for a whole day then take the can off and repeat the test. The can isn't clamped down so the insulation isn't perfect, so I'd expect to see some noticeable fluctuations even with the can on. I've also labeled the long cable for the temperatuse sensor readout (attachments 2 and 3). There will also be an out of loop sensor added in later, but for this test since I am not running the loop it doesn't matter which sensor I monitor. Attachment 4 is a picture of the current setup.

  13858   Thu May 17 13:51:35 2018 Jon RichardsonConfigurationElectronicsDocumentation & Schematics for AUX-PSL PLL

[Jon, Gautam]

Attached is supporting documentation for the AUX-PSL PLL electronics installed in the lower PSL shelf, as referenced in #13845.

Some initial loop measurements by Gautam and Koji (#13848) compare the performance of the LB1005 vs. an SR560 as the controller, and find the LB1005 to be advantageous (a higher UGF and phase margin). I have some additional measurements which I'll post separately.

Loop Design

Pickoffs of the AUX and PSL beams are routed onto a broadband-sensitive New Focus 1811 PD. The AUX laser temperature is tuned to place the optical beat note of the two fields near 50 MHz. The RF beat note is sensed by the AC-coupled PD channel, amplified, and mixed-down with a 50 MHz RF source to obtain a DC error signal. The down-converted term is isolated via a 1.9-MHz low-pass filter in parallel with a 50 Ohm resistor and fed into a Newport LB1005 proportional-integral (PI) servo controller. Controller settings are documented in the below schematic. The resulting control signal is fed back into the fast PZT actuator input of the AUX laser.

Schematic diagram of the PLL.

 

 

 

 

 

 

 

 

 

 

Hardware Photos

Optical layout on the PSL table.

 

PLL electronics installed in the lower PSL shelf.

 

Close-up view of the phase detector electronics.

 

Slow temp. (left) and fast PZT signals into the AUX controller.

 

AUX-PSL beat note locked at 50 MHz offset, from the control room.

 

  13852   Thu May 17 11:56:37 2018 gautamUpdateGeneralEPICS process died on c1ioo

The EPICS process on the c1ioo front end had died mysteriously. As a result, MC autolocker wasn't working, since the autolocker control variables are EPICS channels defined in the c1ioo model. I restarted the model, and now MCautolocker works.

  13851   Thu May 17 09:14:38 2018 SteveUpdateGeneralStack measurement setup decommissioned

The final set-up of stack measurment with 3 load cells and 4 leveling wedge mounts as Atm 1

Sensor voltages BEFORE and AFTER this attempt.

  13850   Wed May 16 21:47:17 2018 KevinUpdatePEMSeismometer Noise Spectra

Earlier today Kira and I reconnected the EX seismometer. I just took some spectra of all three seismometers, shown in the attachments, to compare with past data and to do a rough check of the calibration.

This elog has a spectra from 2010 (GUR1 is now EY) and this elog has one for BS at lower frequencies from 2017. Note that the EX seismometers now have strong peaks that are not at 60 Hz harmonics. Other than these peaks, these old spectra roughly match up with the ones taken today, so the callibration is still roughly the same. I couldn't find any old data for EX (GUR2) though so I don't know for sure that these peaks weren't there before.


gautam 20180517 0930: In 2017, Gur2 (now EX) looked like this. Still peaky, but the peaks seem shifted in frequency. Steve also informed me that the Gur1 and Gur2 cables were swapped n times, so perhaps we shouldn't read too much into that.

  13849   Wed May 16 21:02:22 2018 KevinUpdatePEMADC common mode rejection with new seismometer connections

As described in this elog, the ADC for the seismometers now has the signals wired directly to the ADC instead of going through an AA board or other circuit to remove any common mode noise. This elog describes one test of the common mode rejection of this setup. Guantanamo suggested comparing directly with a recent spectrum taken a few months before the new setup described in this elog.

Today I took a spectrum (attachment 1) of C1:PEM-MIC_2 (Ch17) and C1:PEM-MIC_3 (Ch18) with input to the ADC terminated with 50 Ohms. These are two of the channels plotted in the previous spectrum, though I don't know how that plot was normalized. It's clear that there are now strong 60 Hz harmonic peaks that were not there before, so this new setup does have worse common mode rejection.

  13848   Wed May 16 18:52:50 2018 gautamConfigurationElectronicsPLL mysteries solved

[Koji, Gautam]

Summary:

As I suspected, when the SR560 is operated in 1 Hz, first order LPF mode, the (electronic) transfer function has a zero at ~5kHz (!!!).

Details:

This is what allowed the PLL to be locked with this setting with UGF of ~30kHz. On the evidence of Attachment #3, there is also some flattening of the electrical TF at low frequencies when the SR560 is driving the NPRO PZT. I'm pretty sure the flattening is not a data download error but since this issue needs further investigation anyway, I'm not reading too much into it. I fit the model with LISO but since we don't have low frequency (~1Hz) data, the fit isn't great, so I'm excluding it from the plots.

We also did some PLL loop characterization. We decided that the higher output range (10Vp bs 10Vpp for the SR560) of the LB1005 controller means it is a better option for the PLL. The lock state can also be triggered remotely. It was locked with UGF ~ 60kHz, PM ~45deg.

We also measured the actuation coefficient of the NPRO laser PZT to be 4.89 +/- 0.02 MHz/V. Quoted error is (1-sigma) from the fit of the linear part of the measured transfer function to a single pole at DC with unknown gain. I used the "clean" part of the measurement that extends to lower frequencies for the fit, as can be seen from the residuals plot. Good to know that even though the LDs are dying, the PZT is still going strong :D.

Remaining loop characterization (i.e. verification of correct scaling of in loop suppression with loop gain etc.) is left to Jon.

Measurement schemes:

  1. OLG (Attachment #1) was measured using the usual IN1/IN2 technique.
  2. PZT calibration (Attachment #2) was measured by injecting an excitation at the PLL control point.
    • The ratio of the PLL error point (Volts) to Excitation (Volts) was measured using the SR785.
    • The error point was calibrated by looking at the PLL open loop Vpp (corresponds to pi radians of phase shift).
    • Dividing the fitted gain of the phase->Frequency conversion by the error point calibration, we get the PZT actuation coefficient. 

Some other remarks:

  1. In the swept-sine mode, the SR785 measures transfer functions by taking the ratio of complex FFT values of its inputs at the drive frequency. So the phase in particular is a good indicator of whether the measurement is coherent or not.
  2. In all these measurements, the PLL gain is huge at low frequencies, and hence, the excitation is completely squished on propagating through the loop. E.g. a 10mV excitation is suppressed by a factor of ~60dB = 1000 to 10uV, and if the analyzer autoRange is set to UpOnly, it is easy to see how this is drowned at the IN1 input. This is why the measurements lose coherence below ~1 kHz.
  3. It is easy to imagine implementing an EPICS servo that offloads the DC part of the LB box control signal to the SLOW frequency input on the Lightwave controller. This would presumably allow us to extend the lock timescales. We can also easily implement a PLL autolocker.
  4. Preliminary investigation of the SR560 situation suggests that individual filter stages can only achieve a maximum stopband attenuation of 60dB relative to the passband. When we cascade two stages together, 120dB seems possible...
  13847   Tue May 15 22:11:38 2018 gautamUpdateGeneralIFO maintenance

Since there have been various software/hardware activity going on (stack weighing, AUX laser PLL, computing timing errors etc etc), I decided to do a check on the state of the IFO.

  • c1susaux, c1aux and c1iscaux crates were keyed as they were un-telnet-able.
  • Single arm locking worked fine, TT alignment was tweaked (as these had drifted due to the ADC failure in c1lsc) to maximize Y arm transmission using the dither servos.
  • Arms weren't staying locked for extended periods of time. I particularly suspected ITMX, as I saw what I judged to be excess motion on the Oplev.
  • @Steve - ITMX and BS HeNes look like they are in need of replacement judging by the RIN (although the trend data doesn't show any precipitous drop in power). If we are replacing the BS/PRM Oplev HeNe, might be a good time to plan the inejction path a bit better on that table.
  • RIN in Attachment #1 has been normalized by the mean value of the OL sum channel. There is now a script to make this kind of plot from NDS in the scripts directory (as I found it confusing to apply different calibrations to individual traces in DTT).
  13846   Tue May 15 21:56:57 2018 gautamUpdateGeneralStack measurement setup decommissioned

[steve,koji,gautam]

Since we think we already know the stack mass to ~25% (i.e. 5000 +/- 1000 lbs), we decided to restore the ETMX stack. Procedure followed was:

  • Take photos of all dial indicators and spirit level. We were at ~-22 mils on all 3 indicators, with 0 being the level before we touched the stack two Fridays ago, i.e. May4.
  • Raised all four jacks installed underneat blue crossbeams in 5mil increments until we were at +25mils on all of them. At this point, there was negligible load on the load cells on top of the STACIS legs, and we could easily slide the load cells out.
  • Rotated all jack screws clockwise (i.e. moving jack screws downwards) by 270 degrees. The southeast jackscrew was rotated by an additional 360 degrees. This was to undo all the jack-screw raising we did on Friday, May 4.
  • Re-installed jacks which were present originally on the STACIS legs, taking care to center the jack as best as we could by eye on the STACIS leg, per Dennis Coyne's suggestion not to impose shear strain on STACIS legs. There were supposedly never carrying any load, and are according to Steve, are there more for safety purposes.
  • Lowered all four jacks in 5 mil steps until dial indicators read ~0. The Northwest jack resting on the STACIS leg was somehow ~0.5cm (!!) below the blue crossbeam even though the corresponding dial gauge read 0, so we raised the jack until it was barely grazing the bottom of the blue crossbeam (confirmed by looking at the point where the dial indicator started going up again). Not sure why this should have been, best hypothesis we have is that someone (one of us) changed the level of this jack while it was removed from the setup.
  • Checked that jack screws could not be turned by hand. At this point, all the load has to be resting on the jack screws, as the jacks we had installed to raise the blue crossbeams could be slid out from underneath the blue beams and hence were carrying no load.
  • Took photographs of all dial indicators, spirit level. We were satisfied that we had recovered the "nominal" stack alignment as best as we could judge with the available indicators.
  • ETMX Oplev spot had returned to the PD. ETMX watchdog was re-engaged, optic was re-aligned using SLOW bias sliders to center Oplev spot.
  • EX NPRO was turned back on, and the green beam was readily locked to a cavity TEM00 mode yes.

I will upload the photos to the PICASA page and post the link here later.

Quote:
 

In this case, we only need a mass estimate of the end chamber contents with an accuracy of ~25%. If we think we have that already, we don't need to keep doing the jacks-strain gauge adventure.

 

  13845   Tue May 15 20:51:27 2018 gautamConfigurationElectronicsMaking PLL setup more permanent

[jon, steve, gautam]

Some points which Jon will elaborate upon (and put photos of) in his detailed elog about this setup:

  • PLL electronics (mixer, coupler, ZFL500HLN amplifier and DC power supply for the beatnote, SR560 servo) all reside on the newly installed lower level PSL shelf.
  • Cross connect channel C1:PSL-126MOPA_126CURADJ hijacked for remote temperature control of the AUX NPRO. Note that shield of front panel BNC is ground and so even though the manual says the controller accepts +/-10V, this is not a differential input. BNC cable was routed from cross connect to PSL enclosure, MEDM slider will be setup.
  • There was an SMA cable running from the VEA to the control room which we decided to use for monitoring of the beatnote amplitude on the control room analyzer. Yesterday, Steve and I routed the end of this inside the VEA, near 1X2 originally, to inside the PSL table where it is hooked up to the (20dB) coupled amplifier output. This required some work on the cable tray, we were careful but in case there is some wonkiness in some signals, perhaps this work is to blame.

We are now in a state where the PLL can be locked remotely from the control room by tweaking the AUX laser temperature laugh. Tomorrow, Keerthana will work on getting Craig's/Johannes' Digital Frequency Counter script working here, I think we can easily implement a PLL autolocker if we have some diagostic that tells us if the PLL us locked or not.


Steve informed me that there is an acoustic hum inside the PSL enclosure which wasn't there before. Indeed, it is at ~295Hz, and is from the Bench power supply used to power the ZFL500HLN amplifier. This will have to go...

  13844   Tue May 15 15:13:23 2018 KiraSummaryPEMAcromag issues

I tried calibrating the other channels today, but they still fluctuate. Sometimes they do stabilize at +/- 10V, but then suddenly drop to 5 or 6 V before climbing back up to 10. Turning the legacy off made it go only up to 6.67V. This happens for all the channels, even after doing a factory reset and recalibrating. Not sure what's happening here.

  13843   Tue May 15 13:34:30 2018 SteveUpdatesafetysurf safety training

Pooja and Keirthana received 40m specific basic safety training.

  13842   Tue May 15 10:42:14 2018 KojiUpdatePSLModulation depth measurement for the 3IFO aLIGO EOM and aftermath

The marconi RF output was turned on and thus the RF generator condition was restored to the nominal state on Friday 11th.

  13841   Mon May 14 18:58:32 2018 KevinUpdatePonderSqueezeSqueezing with no SRM
Quote:

Note that for Signal Recycling, which is what Kevin tells us we need to do, there is a DARM pole at ~150 Hz.

To be quantitative, since we are looking at smaller squeezing levels and considering the possibility of using 5 W input power, it is possible to see a small amount of squeezing below vacuum with no SRM.

Attachment 1 shows the amount of squeezing below vacuum obtainable as a function of homodyne angle with no SRM and 5 W incident on the back of PRM. The optimum homodyne angle at 210 Hz is 89.2 deg which gives -0.38 dBvac of squeezing. Figure 2 is the displacement noise at this optimal homodyne angle and attachment 3 is the same noise budget shown as the ratio of the various noise sources to the unsqueezed vacuum.

The other parameters used for these calculations are:

  • 4.5 kΩ series resistance for the ETM coils; 15 kΩ for the ITM coils
  • 100 ppm transmissivity on the folding mirrors giving a PRC gain of 40
  • PD quantum efficiency of 0.88

So maybe it's worth considering going for less squeezing with no SRM if that makes it technically more feasible.

  13840   Mon May 14 08:55:40 2018 Dennis CoyneHowToSEIpreparation of load cell measurement at ETMX

follow up email from Dennis 5-13-2018. The last line agrees with the numbers in elog13821.

Hi Steve & Gautam,

I've made some measurements of the spare (damaged) 40m bellows. Unfortunately neither of our coordinate measurement arms are currently set up (and I couldn't find an appropriate micrometer or caliper), so I could not (yet) directly measure the thickness. However from the other dimensional measurements, and a measurement of the axial stiffness (100 lb/in), and calculations (from the Standards of the Expansion Joint Manufacturers Association (EJMA), 6th ed., 1993) I infer a thickness of 0.010 inch in . This is close to a value of 0.012 in used by MDC Vacuum for bellows of about this size.

I calculate that the maximum allowable torsional rotation is 1.3 mrad. This corresponds to a differential height, across the 32 in span between support points, of 0.041 in.

In addition using the EJMA formulas I find that one can laterally displace the bellows by 0.50 inch (assuming a simultaneous axial displacement of 0.25 inch, but no torsion), but no more than ~200 times. I might be good to stay well below this limit, say no more than ~0.25 inch (6 mm).

If interested I've uploaded my calculations as a file associated with the bellows drawing at D990577-A/v1.

BTW in some notes that I was given (by either Larry Jones or Alan Weinstein) related to the 40m Stacis units, I see a sketch from Steve dated 3/2000 faxed to TMC which indicates 1200 lbs on each of two Stacis units and 2400 on the third Stacis.

  13839   Sun May 13 20:48:38 2018 johannesUpdateGeneralCDS crash

I think the root of the problem is that the /opt/rtapps/ and /cvs/cds/rtapps/ mounting locations point to the same directory on the nfs server. Gautam and I were cleaning up the /cvs/cds/caltech/target/ directory, placing the previous contents of /cvs/cds/caltech/target/c1auxex/, including database files and startup instructions in /cvs/cds/caltech/target/c1auxex_oldVME/, and then moved /cvs/cds/caltech/target/c1auxex2/, which has the channel database and initialization files for the Acromac DAQ, to /cvs/cds/caltech/target/c1auxex/.

This also required updating the systemd entries on c1auxex to point to the changed directory. While confirming that everything worked as before we noticed that upon startup the EPICS IOC complains about not being able to find the caRepeater binary. This was not new and has not limited DAQ functionality in the past, but we wanted to fix this, as it seemed to be some simple PATH issue. While the paths are all correctly defined in the user login shell, systemd runs on a lower level and doesn't know about them. One thing we tried was to let systemd execute /cvs/cds/rtapps/epics-3.14.12.2_long/etc/epics-user-env.sh initializing EPICS. It was strange that the content of that file was pointing to /opt/rtapps/epics-3.14.12.2_long/base, which is not mounted on the slow machines, so we changed the /opt/ it to /cvs/cds/, not realizing that the frontends read from the same directory (as Gautam said, /cvs/cds does not exist as a mount point on the frontend). It ended up not working this way, and apparently I forgot to change it back during clean up. But worse, never elogged it!

In the end, we managed to to give systemd the correct path definitions by explicitly calling them out in /cvs/cds/caltech/target/c1auxex/ETMXenv, to which a reference was added in the systemd service file. The caRepeater warning no longer appears.

  13838   Sun May 13 17:31:51 2018 gautamUpdateGeneralCDS crash

As suspected, this was indeed a path problem. Johannes will elog about it later, but in short, it is related to some path variables being changed in order to try and streamline the EPICS processes on the new c1auxex machine (Acromag Era). It is confusing that futzing around with the slow computing system messes with the realtime system as well - aren't these supposed to be decoupled? Once the paths were restored by Johannes, everything compiled and restarted fine. We even have a beam on the AS camera, which was what triggered this whole thingyes.

Anyways, Attachment #1 shows the current status. I am puzzled by the red TIMING indicators on the c1x04 and c1x02 processes, it is absent from any other processes. How can this be debugged further?

Quote:
 

I suspect this is some kind of path problem - the EPICS_BASE bash variable is set to /cvs/cds/rtapps/epics-3.14.12.2_long/base on the FEs, while /cvs isn't even mounted on the FEs (nor do I think it should be). I think the correct path should be /opt/rtapps/epics-3.14.12.2_long/base. Why should this have changed?

  13837   Sun May 13 15:15:18 2018 gautamUpdateGeneralCDS crash

I found the c1lsc machine to be completely unresponsive today. Looking at the trend of the state word, it happened sometime yesterday (Saturday). The usual reboot procedure did not work - I am not able to bring back any of the models on any of the machines, during the restart procedure, they all fail. The logfile reads (for the c1ioo front end, but they all behave the same):

[  309.783460] c1x03: Initializing space for daqLib buffers
[  309.887357] CPU 2 is now offline
[  309.887422] c1x03: Sync source = 4
[  309.887425] c1x03: Waiting for EPICS BURT Restore = 2
[  309.946320] c1x03: Waiting for EPICS BURT 0
[  309.946320] c1x03: BURT Restore Complete
[  309.946320] c1x03: Corrupted Epics data:  module=0 filter=1 filterType=0 filtSections=134610112
[  309.946320] c1x03: Filter module init failed, exiting
[  363.229086] c1x03: Setting stop_working_threads to 1
[  364.232148] DXH Adapter 0 : BROADCAST - dx_user_mcast_unbind - mcgroupid=0x3
[  364.233689] Will bring back CPU 2
[  365.236674] Booting Node 1 Processor 2 APIC 0x2
[  365.236771] smpboot cpu 2: start_ip = 9a000
[  309.946320] Calibrating delay loop (skipped) already calibrated this CPU
[  365.251060] NMI watchdog enabled, takes one hw-pmu counter.
[  365.252135] Brought the CPU back up
[  365.252138] c1x03: Just before returning from cleanup_module for c1x03

Not sure what is going on here, or what "Corrutped EPICS data" is supposed to mean. Thinking that something was messed up the last time the model was compiled, I tried recompiling the IOP model. But I'm not able to even compile the model, it fails giving the error message

make[1]: Leaving directory '/opt/rtcds/caltech/c1/rtbuild/3.4'
make[1]: /cvs/cds/rtapps/epics-3.14.12.2_long/modules/seq/bin/linux-x86_64/snc: Command not found
make[1]: *** [build/c1x03epics/c1x03.c] Error 127
Makefile:28: recipe for target 'c1x03' failed
make: *** [c1x03] Error 1

I suspect this is some kind of path problem - the EPICS_BASE bash variable is set to /cvs/cds/rtapps/epics-3.14.12.2_long/base on the FEs, while /cvs isn't even mounted on the FEs (nor do I think it should be). I think the correct path should be /opt/rtapps/epics-3.14.12.2_long/base. Why should this have changed?

I've shutdown all watchdogs until this is resolved.

  13836   Sat May 12 10:02:03 2018 ranaUpdateGeneralMore refinement of DARM control signal projection

Good question! I've never calculated what the resonance frequency would be if had an inductive resistor with our cable capacitance (~50 pF/m I guess).

ELOG V3.1.3-